OpenStudy (anonymous):

the sum of two consecutive integers is 159...what are they?

OpenStudy (anonymous):

x+(x+1)=159

OpenStudy (anonymous):

79 and 80

OpenStudy (anonymous):

2x+1=159 2x=158 x=79 and the other number is 80

OpenStudy (anonymous):

Let two conscutive no will be....x and x+1 so x+x+1=159 2x=158 x=79 now we have x=79 so other no will be x+1=79+1=80